Science and medicine. A passenger train can travel 325 mi in the same time a freight train takes to travel 200 mi. If the speed of the passenger train is 25 mi/h faster than the speed of the freight train, find the speed of each.
Answer by checkley71(8403)

325/(x+25)=200/x now cross multiply
325x=200(x+25)
325x=200x+5000
325x-200x=5000
125x=5000
x=5000/125
x=40 mph is the speed of the freight train.
40+25=65 mph is the speed of the pasenger train.
proof
325/65=200/40
5=5